# Sparse is Enough in Fine-tuning Pre-trained Large Language Model

### v1

@@ 3 Understanding Pre-training-fine-tuning from a Gradient-based Perspective
3.1 Visualization of Loss Landscape
Loss landscape is a high-dimensional geometry which is completely determined as soon as a dateset and the network structure are specified(Li et al., 2018a).For high-dimensional geometries such as loss landscapes, it is not possible to imagine and understand their geometric properties as intuitively as in low-dimensional shapes. Despite the difficulty in understanding such high-dimensional geometries, researchers have proposed methods to project high-dimensional loss surfaces into low-dimensional space, including one-dimensional visualization and two-dimensional visualization, in order to easily understand and study some of the properties of loss landscapes. (Hao et al., 2019) first attempts to explain the role of pre-training and fine-tuning in BERT from the perspective of the loss landscape, which visualizes the optimization trajectories both of random initialization and pre-trained initialization in one and two dimensions respectively, giving an intuitive explanation for the faster convergence and better generalization under the pre-training fine-tuning paradigm. However, they treat training from scratch and fine-tuning from pre-trained as two separated cases, ignoring the transition of loss landscape from random initialization to pre-trained initialization. Next, we visualize and explore the transition of loss landscape.
In exploring the transition of the loss landscape from random initialization to pre-trained initialization, we note $\theta_{0}$ as the randomly initialized parameters and $\theta_{1}$ as the pre-trained parameters. (Goodfellow et al., 2015) plots the loss landscape curve from $\theta_{0}$ to $\theta_{1}$ by linearly sampling and interpolating the network parameters between $\theta_{0}$ and $\theta_{1}$, the loss curve $f$ is defined as:
$f(\alpha)$ $=\mathcal{L}((1-\alpha)\theta_{0}+\alpha\theta_{1})$
$=\mathcal{L}(\theta_{0}+\alpha(\theta_{1}-\theta_{0}))$
$=\mathcal{L}(\theta_{0}+\alpha\delta)$
where $\alpha\in[0,1]$ and $\delta$ represents a direction.
The two-dimensional loss surface $f$ is plotted similarly to the one-dimensional one, determined by a starting point and two directions:
In our setting, $\delta_{1}=\theta_{1}-\theta_{0},\alpha\in[-0.5,1.5],\beta\in[-1,1]$, when $\alpha=0,\beta=0$, the model is in the random initialization state, and when $\alpha=1,\beta=0$, the model is in the pre-trained state. We use a method similar to (Li et al., 2018b) to construct the direction $\delta_{2}$ to avoid the scaling effect between different components, multiplying $\delta_{1}$ by a Gaussian random number points by points, i.e. $\delta_{2_{i}}=d_{i}*\delta_{1_{i}},$ where $d_{i}$ follows a normal distribution.
We investigate the transition of the loss landscape using RoBERTa-base(Liu et al., 2019) on a range of tasks at GLUE, some of which are shown in the figure 1. The full results and specific experimental details are given in the Appendix.
A prior idea would be that pre-trained models are more efficient at adapting to downstream tasks than training from scratch because the knowledge learned during pre-training reduces the initial loss. However, a striking fact is that pre-training does not make all tasks have lower losses than random initialization at the beginning of fine-tuning, e.g. MNLI, MRPC, etc. Instead, for most task, there are extremely higher losses than random initialization around the pre-trained initialization. A common feature of the loss landscape across tasks is that the oscillations of the loss around random initialization are generally small in magnitude, whereas the loss oscillations around pre-training are sharp, i.e. after a valid pre-training process, the loss landscape shifts from low amplitude oscillations to high amplitude oscillations. Since the 1-D and 2-D loss landscapes are projections of the original loss landscape, we are curious whether the special properties exhibited in these low-dimensional projections reflect a general property of the high-dimensional space in a statistical sense, that is, whether there is some consistency across most dimensions.
@@ 3.2 Quasi-Sparse Gradient Distribution
(Ye et al., 2020; Wiedemann et al., 2020) take the assumption that the distribution of the parameters’ gradients approximates a normal distribution. In the figure 2, we plot the gradient distribution of one parameter in RoBERTa-Large on the MNLI dataset. Both the gradient of the model trained from scratch and from pre-trained show a bell-shaped distribution with a mean of zero. Under the same histogram settings, the gradient distributions of from scratch and pre-trained models show a huge difference in frequency and value range. Figure 2(c) displays the pre-trained gradient distribution on a log-scale, where the majority of gradients are concentrated around zero, with only a very small portion having larger gradient values, and the maximum value of pre-trained is about two orders of magnitude higher than that from scratch. Figure 2(d) shows the gradient proportion of the components with the top xabsolute gradient value. The pre-trained gradient shows a more extreme property, where 1% of the components account for 99% of the total gradient norm. This severe imbalance in distribution exhibits a property similar to sparsity, which we refer to as Quasi-Sparse. The majority of information in the gradient space can be represented by a very small number of components, leading to a low intrinsic dimension of the objective loss landscape(Li et al., 2018a). This imbalance in distribution also explains the transition from low to high amplitude oscillation discussed in 3.1. With the same step size, a balanced distribution of gradients in each dimension would lead a more smooth change in loss, whereas in the case of an imbalanced distribution of gradients, the dimensions with extreme gradients will dominate the change in loss, resulting in strong oscillations.
3.3 Flatter Minima
A popular view is that flatter minima will generalize better(Hochreiter & Schmidhuber, 1997; Jiang et al., 2019; Liu et al., 2022), and for pre-trained models, even with the same pre-training loss, flatter minima have better downstream performance(Liu et al., 2022). (Foret et al., 2021; Bahri et al., 2022) introduce sharpness-aware regularization to reduce the sharpness of the minima during training as a way to improve the generalization ability of the model. For the pre-trained model on the downstream task, whose gradient distribution is akin to sparsity, finding flat minima becomes less difficult, with the flat dimensions dominating the landscape, leaving the sharp dimensions to account for only a very small portion. By stochastic gradient descent methods, it is easy to search for flat minima in a relatively small subspace spanned by the sharp direction, i.e., a minimizer that can sufficiently generalize.
@@ 4.1 SIFT: Sparse Increment Fine-Tuning
We define the fine-tuned parameters as the sum of the pre-trained parameters and an increment, $x_{ft}=x_{pt}+\Delta x.$ Updating only a portion of the components means that $\Delta x$ is a sparse matrix, hence we name this update method Sparse Increment Fine-Tuning (SIFT). Below, we will explain that due to the special gradient properties of the pre-trained model, $\Delta x$ can ensure sufficient descent of the loss.
The idea of only updating partial components of the parameters is similar to the subspace method for non-linear optimization (Liu et al., 2021). In SIFT, we aim to update only the components with the top xabsolute values of the gradient, so the update direction will be in a subspace spanned by these components’ dimensions. We adopt the same definition as in (Liu et al., 2021), where $g_{k}^{i}$ is denoted as the $i^{th}$ component of the gradient $g_{k}$. Sorted in descending order of absolute value, the top xcomponents are selected, that is:
@@
Therefore, the update direction $d$ are in the subspace $\mathcal{S}=\text{span}\{e^{i_{1}},e^{i_{2}},\ldots,e^{i_{n}}\}$, called the $\tau$-steepest coordinates subspace, where $e_{i}$ refers to the $i^{th}$ standard basis vector. Based on this, we define the steepest descent direction as sufficiently descending, when,
@@
Figure 3: (Left) compares the difference in the proportion of gradients accounted for by two different top 1% component selection methods: the top 1% of the current batch and the top 1% of the first batch. Although sticking with the first batch’s top 1% as opposed to selecting the top 1% of the current batch results in some loss of gradient information, the difference remains within an acceptable range. Avoiding frequent changes of components can improve training efficiency and preserve the historical information of Adam-like optimizers. In contrast, (Right) shows the gradient proportion variation in the model trained from scratch, where fixing the top 1% of the first batch leads to a greater loss of the gradient information.
Due to the limitation of computational cost, we usually cannot obtain the gradients over the complete data. Instead, stochastic gradient type methods typically take a mini-batch sampled from the full data. Without the complete gradient information, SIFT uses the gradient from the first batch(or first several batches by using gradient accumulation) as an estimate of the complete gradient and selects the components to be updated based on this. A natural question arises: do the components selected based on the gradient of the first batch still effectively represent the full gradient information in subsequent batches? That is, whether there will be significant differences between the gradients of different batches. As shown in Figure 3, we compare the proportion of the top 1% components of each batch’s gradients with the proportion of the top 1% components of the first batch’s gradients. For pre-trained models, using the top xcomponents of the first batch does indeed result in some loss of gradient information, but except for some outlier samples, the difference always remains within an acceptable level compared to the top 1% of the current batch. In contrast, the top 1% components of the first batch in models trained from scratch tend to lose more gradient information. Therefore, for efficiency, SIFT always updates the top xcomponents of the first batch in the whole training process. This approach enables the use of Adam-like optimizers that require historical state information, as frequent changes of components would result in the loss of their historical information. Appropriate periodic changes of the components that need to be updated is also an acceptable choice, and this could be one of the focus for future work.
@@ 4.2 A Memory-efficient Implementation of SIFT
The key to SIFT lies in obtaining the gradients of partial components. Under existing deep learning frameworks (such as PyTorch(Paszke et al., 2017)), gradients of all parameters are able to be accessed only after backward propagation. A simple implementation of SIFT is to retrieve the gradients of partial components through indexing after all gradients have been calculated. However, as all gradients have already been stored, storing the gradients of partial components at this point does not reduce the memory overhead of gradient storage. The memory overhead during training mainly consists of four parts: parameters, gradients, optimizer states, and activations. Although this implementation does not reduce the storage overhead of gradients, it can significantly reduce the storage overhead of the optimizer states.
(Lv et al., 2023) proposes an alternative SGD method that reduces gradient storage by inserting a hook function, thereby fusing the gradient into the parameters as soon as the gradient is computed. Inspired by (Lv et al., 2023), we propose a memory-efficient implementation of SIFT. As shown in Figure 4, we register a Sparse Parameter (SP) and corresponding Sparse Gradient (SG) for the parameters to be updated, storing them in a manner that preserves values and indexes. SP does not obtain Sparse Gradient SG through normal backward propagation. Instead, we inject a hook function into the parameters to capture the computed gradients and assign them to SG via indexing. The Sparse Parameter SP with Sparse Gradient SG is updated through the normal optimizer, and finally, SP is integrated into the initial parameter P.
GC Params Grad Optim States Activation Total
Full 12.55 12.55 50.21 35.72 111.03
0.81 76.12
SIFT 12.55 0.626 2.51 35.72 51.41
0.81 16.50
Through this method, for xsparse updates, we can simultaneously reduce the gradients and optimizer states to the original x%. Combined with techniques such as mixed-precision training and gradient checkpointing, it is possible to fine-tune a 7B model on a single RTX 3090 24GB.
@@ 5 Experiments
Our evaluation of SIFT mainly consists of two parts, targeting two use cases: Natural Language Understanding (NLU) and Natural Language Generating (NLG). For NLU, we choose to align with the evaluation on LoRA (Hu et al., 2021), selecting the GLUE Benchmark (Wang et al., 2019) as our evaluation dataset, and compare it with the data reported in (Hu et al., 2021). For NLG, to validate the effectiveness of SIFT on much larger models, we choose instruction-tuning as our test case. We select Llama (Touvron et al., 2023) as our backbone models, use the alpaca dataset (Taori et al., 2023) for instruction-tuning, and conduct evaluations on datasets such as MMLU (Hendrycks et al., 2021) and HumanEval (Chen et al., 2021). In the experiment, similar to (Hu et al., 2021), we only consider updating the parameters in the self-attention modules. We further analyze the choice of parameters to update in section 6. By adjusting the sparsity rate of SIFT, we ensure a consistent number of trainable parameters.
In section 5.3, we conduct further experimental analysis towards SIFT. Since SIFT selects components to update through gradient analysis, we compare SIFT with random component selection in section 5.3.2. Additionally, as we heuristically choose to perform sparse updates on q,k,v and o, we compare the specific performance of applying SIFT in different parameters. Finally, we explore the impact of sparsity rate on the downstream task, investigating what is the limit of sparsity rate required for an acceptable performance in the downstream task.
The evaluation includes the following baselines:
  • Full fine-tune: All pre-trained parameters are updated.
  • Adapter: (Houlsby et al., 2019) proposes to adapt pre-trained models to downstream tasks by inserting trainable Adapter modules into the self-attention layers while keeping the pre-trained parameters frozen. These modules have a bottleneck structure with non-linear function. We denote this original version as $\text{Adpt}^{H}$. Based on (Houlsby et al., 2019), (Pfeiffer et al., 2021) proposed a more efficient version of the Adapter modules, referred to as $\text{Adpt}^{P}$.
  • LoRA(Hu et al., 2021): freezes the weights of the pre-trained model and adds trainable rank decomposition matrices in parallel to each layer of the Transformer architecture.
@@ 5.1 GLUE Benchmark
We evaluate SIFT on the GLUE benchmark (Wang et al., 2019). We select RoBERTa (Liu et al., 2019) as the backbone model for testing. The general experimental setup is consistent with (Hu et al., 2021) and we apply a 0.8% sparsity rate with SIFT to the Query, Key, Value, and Output projection in the self-attention module to ensure the number of trainable parameters is consistent with the compared baselines. See Appendix for detailed hyperparameter settings. Table 2 reports the results of pre-trained Roberta-large with different fine-tuning methods on the GLUE benchmark and the number of trainable parameters (excluding the classifier). The results of each run are taken from the best epoch. Table 2 reveals that SIFT performs competitive on the GLUE benchmark compared to full fine-tuning and other PEFT methods.

### v3

@@ 3.2 Visualization of Loss Landscape
If the probability distributions of gradients on each dimension $e_{i}$ are relatively similar, meaning there is an expectation $E(\nabla L_{e_{i}}(\theta_{0}))\approx E_{0}$, where $E_{0}$ is a constant, and since $N$ is usually quite large, the loss can be further approximated as $L(\theta_{0})+\alpha\sum d_{i}E_{0}$. This indicates that the loss will not exhibit sharp oscillations. However, if extreme distributions exist in certain dimensions, for example, if the $N^{th}$ dimension has a different distribution, then the loss can be approximated as $L(\theta_{0})+\alpha\sum_{i=1}^{N-1}d_{i}E_{0}+\alpha d_{N}\nabla L_{e_{N}}(\theta_{0})$. In this case, the loss will be significantly influenced by the gradient of the $N^{th}$ dimension. Therefore, a flat loss landscape implies similar distributions among dimensions and the intense oscillations in loss landscape indicate that certain dimensions dominate the changes of the loss, suggesting an inconsistency in the distributions among dimensions. We verify this inconsistency in gradient distributions in the following Section 3.3 and highlight the quasi-sparsity in gradients of pre-trained models.
@@ 3.3 Quasi-Sparse Gradient Distribution
Ye et al. (2020); Wiedemann et al. (2020) take the assumption that the distribution of the parameter gradients approximates a normal distribution. In Figure 3, we plot the gradient distribution of one parameter in RoBERTa-Large on the MNLI dataset and other parameters also have the similar property shown in the following part. Both the gradients of the model trained from scratch and from pre-trained show a bell-shaped distribution with a mean of near zero. Under the same histogram settings, the gradient distributions of from scratch and pre-trained models show a huge difference in frequency and value range. Figure 3(c) displays the pre-trained gradient distribution on a log-scale, where the majority of gradients are concentrated around zero, with only a very small portion having larger gradient values, and the maximum value of pre-trained is about two orders of magnitude higher than that from scratch. Figure 3(d) shows the gradient proportion of the components with the top xabsolute gradient value. The pre-trained gradient shows a more extreme property, where 1% of the components account for 99% of the total gradient norm. This severe imbalance in distribution exhibits a property similar to sparsity, which we refer to as Quasi-Sparse. The majority of information in the gradient space can be represented by a very small number of components.
Through pre-training, the model reaches a sub-manifold in the parameter space where the parameters are with higher distribution probabilities and can be expressive of language (corresponding to the bell-shaped distribution in Figure 1). Therefore, during the fine-tuning process on downstream tasks, it is only necessary to search within this manifold to find a good solution. This results in redundancy on most dimensions, manifested by smaller gradient values. Therefore, the compressed searching space is the key to the low intrinsic dimension of the objective loss landscape (Li et al., 2018a). A natural question is whether updating only in these dimensions can effectively adapt the pre-trained model to downstream domains. A prior idea is that, due to the quasi-sparsity of the gradient, updating only partial components with larger gradients can also effectively reduce the empirical loss, while the generalization loss can also be consistently reduced due to tighter upper bounds on the generalization error of pre-trained models.
In the following Section 4, we propose a memory-efficient implementation of the sparse fine-tuning scheme, which validates the feasibility of fine-tuning language models entirely in a component-sparse way.
@@ 4.1 SIFT: Sparse Increment Fine-Tuning
We define the fine-tuned parameters as the sum of pre-trained parameters and an increment, $x_{ft}=x_{pt}+\Delta x.$ Updating only a portion of the components means that $\Delta x$ is a sparse matrix, hence we name this update method Sparse Increment Fine-Tuning (SIFT). Below, we will explain that due to the quasi-sparse gradient property of the pre-trained model, $\Delta x$ can ensure sufficient descent of the loss.
The idea of only updating partial components of the parameters is similar to the subspace method for non-linear optimization (Liu et al., 2021). In SIFT, we aim to update only components with the top xabsolute values of the gradient, so the update direction will be in a subspace spanned by these components’ dimensions. We adopt the definition in (Liu et al., 2021), where $g_{k}^{i}$ is denoted as the $i^{th}$ component of gradient $g_{k}$. Sorted in descending order of absolute value, the top xcomponents are selected:
@@
Therefore, the update direction $d$ are in the subspace $\mathcal{S}=\text{span}\{e^{i_{1}},e^{i_{2}},\ldots,e^{i_{n}}\}$, called the $\tau$-steepest coordinates subspace, where $e_{i}$ refers to the $i^{th}$ standard basis. Based on this, we define the steepest descent direction as sufficiently descending, when,
@@
Figure 4: (Left) compares the difference in the proportion of gradients accounted for by two different top 1% component selection methods: the top 1% of the current batch and the top 1% of the first batch. Although sticking with the first batch’s top 1% as opposed to selecting the top 1% of the current batch results in some reductions of gradient information, the difference remains within an acceptable range. Avoiding frequent changes of components can improve training efficiency and preserve the historical information of Adam-like optimizers. In contrast, (Right) shows the gradient proportion variation in the model trained from scratch, where fixing the top 1% of the first batch leads to a greater reduction of the gradient information.
Due to the limitation of computational cost, we usually cannot obtain the gradients over the complete data. Instead, stochastic gradient type methods typically take a mini-batch sampled from the full data. Without the complete gradient information, SIFT uses the gradient from the first batch(or first several batches by using gradient accumulation) as an estimate of the complete gradient and selects the components to be updated based on this. A natural question arises: do the components selected based on the gradient of the first batch still effectively represent the full gradient information in subsequent batches? That is, whether there will be significant differences between the gradients of different batches. As shown in Figure 4, we compare the proportion of the top 1% components of each batch’s gradients with the proportion of the top 1% components of the first batch’s gradients. For pre-trained models, using the top xcomponents of the first batch does indeed result in some reduction of gradient information, but except for some outlier samples, the difference always remains within an acceptable level compared to the top 1% of the current batch. In contrast, the top 1% components of the first batch in models trained from scratch tend to lose more gradient information. Therefore, for efficiency, SIFT always updates the top xcomponents of the first batch in the whole training process. This approach enables the use of Adam-like optimizers that require historical state information, as frequent changes of components would result in the loss of their historical information. Appropriate periodic changes of the components that need to be updated is also an acceptable choice, and this could be one of the focus for future work.
@@ 4.2 A Memory-efficient Implementation of SIFT
The key to SIFT lies in obtaining gradients of partial components. Under existing deep learning frameworks (such as PyTorch (Paszke et al., 2017)), gradients of all parameters are able to be accessed only after backward propagation. A simple implementation of SIFT is to retrieve the gradients of partial components through indexing after all gradients have been calculated. However, as all gradients have already been stored, storing the gradients of partial components at this point does not reduce the memory overhead of gradient storage. The memory overhead during training mainly consists of four parts: parameters, gradients, optimizer states, and activation. Although this implementation does not reduce storage overhead of gradients, it can significantly reduce that of the optimizer states.
@@
Lv et al. (2023) proposes an alternative SGD method that reduces gradient storage by inserting a hook function, thereby fusing the gradient into the parameters as soon as the gradient is computed. Inspired by Lv et al. (2023), we propose a memory-efficient implementation of SIFT. As shown in Figure 5, we register a Sparse Parameter (SP) and corresponding Sparse Gradient (SG) for the parameters to be updated, only storing values and indexes for each. SP does not obtain SG through normal backward propagation. Instead, we inject a hook function into the parameters to capture the computed gradients and assign them to SG via indexing. SP with SG is updated through the normal optimizer, and finally, SP is integrated into the initial parameter P.
GC Params Grad Optim States Activation Total
Full 12.55 12.55 50.21 35.72 111.03
0.81 76.12
SIFT 12.55 0.626 2.51 35.72 51.41
0.81 16.50
For xsparse updates, we can simultaneously reduce the gradients and optimizer states to the original x%. Combined with techniques such as mixed-precision training and gradient checkpointing, it is possible to fine-tune a 7B model on a single RTX 3090 24GB, as shown in Table 2.
@@ 5 Experiments
Our evaluation mainly consists of two parts, targeting two cases: Natural Language Understanding (NLU) and Natural Language Generating (NLG). For NLU, we align with the evaluation of LoRA (Hu et al., 2021), selecting the GLUE Benchmark (Wang et al., 2018) as our evaluation dataset, and compare it with the data reported in (Hu et al., 2021). For NLG, to validate the effectiveness of SIFT on much larger models, we choose instruction-tuning as our test case. We adopt Llama (Touvron et al., 2023) as our backbone models, use the alpaca dataset (Taori et al., 2023) for instruction-tuning, and conduct evaluations on benchmarks such as MMLU (Hendrycks et al., 2020) and HumanEval (Chen et al., 2021). In the experiments, similar to (Hu et al., 2021), we only consider updating the parameters in the self-attention modules. By adjusting the sparsity rate of SIFT, we ensure a consistent number of trainable parameters. See Appendix B for detailed hyper-parameter settings, categories scores of MMLU and generated samples of HumanEval.
In Section 5.3, we conduct further experimental analysis towards SIFT. Since SIFT selects components to update through gradient analysis, we compare SIFT with random component selection in Section 5.3.1. Additionally, as we heuristically choose to perform sparse updates on q,k,v and o, we compare the specific performance of applying SIFT in different parameters. Finally, in Section 5.3.2, we explore the impact of sparsity rate on the downstream task, investigating what is the limit of sparsity rate required for an acceptable performance in the downstream task. The evaluation includes the following baselines: Full fine-tune: All pre-trained parameters are updated. Adapter: (Houlsby et al., 2019) proposes to adapt pre-trained models to downstream tasks by inserting trainable Adapter modules into the self-attention layers while keeping the pre-trained parameters frozen. These modules have a bottleneck structure with non-linear function. We denote this original version as $\text{Adpt}^{H}$. Based on (Houlsby et al., 2019), (Pfeiffer et al., 2020) proposed a more efficient version of the Adapter modules, referred to as $\text{Adpt}^{P}$. LoRA(Hu et al., 2021): freezes the weights of the pre-trained model and adds trainable rank decomposition matrices in parallel to each layer of the Transformer architecture.
@@ 5.1 GLUE Benchmark
We evaluate SIFT on the GLUE benchmark (Wang et al., 2018). We select RoBERTa (Liu et al., 2019) as the backbone model for testing. The general experimental setup is consistent with (Hu et al., 2021) and we apply a 0.8% sparsity rate with SIFT to the Query, Key, Value, and Output projection in the self-attention module to ensure the number of trainable parameters is consistent with the compared baselines. Table 1 reports the results of pre-trained RoBERTa-Large with different fine-tuning methods on the GLUE benchmark and the number of trainable parameters (excluding the classifier). The results of each run are taken from the best epoch. Table 1 reveals that SIFT performs competitive on the GLUE benchmark compared to full fine-tuning and other PEFT methods.
5.2 Instruction-tuning
Wei et al. (2021) proposes to perform supervised fine-tuning on top of pre-trained language models to make them align instructions and generate more meaningful content. Taori et al. (2023) fine-tunes pre-trained Llama on a 52k instruction-following dataset (Alpaca), which is generated by the technique introduced by Wang et al. (2022). We fine-tune Llama with Alpaca in different ways including full fine-tuning, LoRA and our SIFT. We evaluate language models on two benchmarks including MMLU and HumanEval. For MMLU, the evaluation metric is the accuracy of the generating answer in zero-shot setting, while HumanEval is evaluated by pass@1 and pass@10 metric. Higher values are better for all metrics. Table 3 demonstrates that sparse structures like SIFT can inject knowledge into pre-trained LLMs and exhibit a competitive performance.